# Molecular Identification and Peak Assignment: Leveraging Multi-Level Multimodal Alignment on NMR

### v1

@@ Appendix B Further ablation study about parameter choices
B.2 Ablation Study on the Choice of $\tau_{1}$ and $\tau_{2}$
We also conducted a further ablation study exploring different combinations of $\tau_{1}$ and $\tau_{2}$ as shown in Table B.2. For this analysis, we fixed the GIN depth at 5, set the GIN embedding dimensionality to 128, and maintained a projection dimension of 512. We observe that the best performance is achieved when $\tau_{1}$ = $10^{-5}$ and $\tau_{2}$ = $10^{1}$.
Table B.2: Ablation study about $tau_{1}$ and $tau_{2}$. We have 5 layers and 128 dimension as the final representation.
@@ B.2 Ablation Study on the Choice of $\tau_{1}$ and $\tau_{2}$
Appendix C Additional Results on Molecular Retrieval
Table C.1: Zero-shot molecular retrieval top 5, 10, 25 accuracy (%) with K-M3AID and baselines
@@ Appendix C Additional Results on Molecular Retrieval
Appendix D Additional Discussion about Isomers
D.1 Isomer Category
@@
Isomers typically fall into two main categories: constitutional (structural) isomers, which share the same chemical formula but display distinct atom connectivity, and stereoisomers (spatial isomers), which share the same topology graph but diverge in their three-dimensional arrangement (see Figure D.1). Constitutional isomers are NMR-variant, meaning that different isomers produce distinct NMR spectrum. In the sub-categories of stereoisomers, enantiomers are NMR-invariant, but diastereomers and cis-trans isomers are NMR-variant.
Figure D.1: NMR Variability in Isomers
D.2 Isomers Group for $\text{C}_{7}\text{H}_{11}\text{NO}_{3}$
@@
Here is an example for isomer groups. In this isomer group of $\text{C}_{7}\text{H}_{11}\text{NO}_{3}$, they all share the same chemical formula in Figure D.2. The first 10 are constitutional (structural) isomers of each other (cycled green), the last 4 are two pairs of diastereomers (cycled brown). Each of these isomers corresponds to a distinct NMR spectrum.
Figure D.2: Isomer demo for $\text{C}_{7}\text{H}_{11}\text{NO}_{3}$
Appendix E Additional Results on Peak Assignment
@@
In complex natural product molecules, it is a common situation that the local contents of some atoms within the same molecule exhibit a high degree of similarity. It gives rise to challenges for the atomic alignment, as some atoms correspond to ppm values in close proximity. However, our K-M3AID model is capable of recognizing each of the atoms with effective learnt embeddings and deciphering the correspondences among the atoms and the peaks at zero-shot. Two complex natural product molecules with multiple rings (4 and 4, respectively) and multiple chiral centers (6 and 8, respectively) are taken to showcase the effectiveness of atomic alignment (see Appendix Figure E.1).
In molecular A in Figure  E.2, atom 13 and atom 14 are tertiary carbons (attaching to 3 carbons and 1 hydrogen) and on the same 5-member ring, corresponding to the ppm of 34.3 and 35.6, respectively. The similar local content of these two atoms fools SP-ID and WP-ID. In addition, WP-ID fails with more atomic alignments. The molecular B is chemical symmetric regarding atom 0. Thus, atom 1 and atom 3 correspond to the same peak on the spectra. The ppm of atom 1 and atom 3 is 114.2, the ppm of atom 2 and atom 4 is 110.0. While there is 4.2 difference, SP-ID and WP-ID fails to pick up right alignment for atom 1 and atom 3. In contrast, K-ID succeed to align the atoms with peaks in both molecules.
@@
Figure E.1: Examples of Zero-shot Atomic Alignment for Complex Natural Products. Yellow cells in the PPM difference represent the ground truth alignment.
@@
Figure E.2: Extra case studies of IE-Meta-MMA.Yellow cells in the PPM differerence represent the ground truth alignment, and red cross represents the wrong alignment.

### v2

@@ Appendix C Further ablation study about parameter choices
C.2 Ablation Study on the Choice of $\tau_{1}$ and $\tau_{2}$
We also conducted a further ablation study exploring different combinations of $\tau_{1}$ and $\tau_{2}$ as shown in Table C.2. For this analysis, we fixed the GIN depth at 5, set the GIN embedding dimensionality to 128, and maintained a projection dimension of 512. We observe that the best performance is achieved when $\tau_{1}$ = $10^{-5}$ and $\tau_{2}$ = $10^{1}$.
Table C.2: Ablation study about $tau_{1}$ and $tau_{2}$. We have 5 layers and 128 dimension as the final representation.
@@ C.2 Ablation Study on the Choice of $\tau_{1}$ and $\tau_{2}$
Appendix D Additional Results on Molecular Retrieval
Table D.1: Zero-shot molecular retrieval top 5, 10, 25 accuracy (%) with K-M3AID and baselines
@@ Appendix D Additional Results on Molecular Retrieval
Table D.2: Comparing sampled molecules to their Top 1 neighbors using K-M3AID, SP-ID, WP-ID (th = 1), and K-M3AID without communication across datasets of varying sizes and employing different similarity metrics (%) including Cosine (Salton & McGill, 1986)), Dice (Dice, 1945)), Russel (Russel, 1980), Sokal (Sokal & Sneath, 1963) and Tanimoto (Tanimoto, 1957).
Methods Similarity Metric $10^{2}$ $10^{3}$ $10^{4}$ $10^{5}$ $10^{6}$
K-M3AID Cosine 96.1$\pm$0.9 81.9$\pm$3.5 50.0$\pm$1.2 23.9$\pm$0.7 13.5$\pm$1.5
Dice 96.1$\pm$0.9 81.9$\pm$3.5 50.0$\pm$1.2 23.7$\pm$0.6 13.1$\pm$1.5
Russel 95.8$\pm$0.1 80.4$\pm$3.9 46.4$\pm$1.1 18.1$\pm$0.8 5.9$\pm$1.7
Sokal 95.9$\pm$0.9 80.8$\pm$3.8 47.2$\pm$1.2 19.5$\pm$0.7 7.7$\pm$1.7
Tanimoto 96.0$\pm$0.9 81.2$\pm$3.7 48.2$\pm$1.2 21.0$\pm$0.7 9.6$\pm$1.6
SP-ID Cosine 95.4$\pm$0.8 78.6$\pm$2.4 42.5$\pm$3.6 20.9$\pm$1.5 12.2$\pm$0.7
Dice 95.5$\pm$0.8 78.6$\pm$2.5 42.4$\pm$3.6 20.3$\pm$1.5 11.8$\pm$0.8
Russel 95.1$\pm$0.8 76.9$\pm$2.7 38.3$\pm$3.8 14.4$\pm$1.6 4.3$\pm$0.9
Sokal 95.2$\pm$0.8 77.4$\pm$2.6 39.4$\pm$3.8 15.9$\pm$1.6 6.3$\pm$0.8
Tanimoto 95.3$\pm$0.8 77.8$\pm$2.6 40.4$\pm$3.7 17.4$\pm$1.6 8.2$\pm$0.8
WP-ID(th=1) Cosine 93.4$\pm$0.6 73.9$\pm$0.1 37.5$\pm$1.2 17.2$\pm$0.4 11.4$\pm$0.5
Dice 93.3$\pm$0.6 73.7$\pm$0.9 37.3$\pm$1.2 16.9$\pm$0.4 11.1$\pm$0.5
Russel 92.9$\pm$0.6 71.7$\pm$1.0 32.7$\pm$1.3 10.8$\pm$0.5 3.7$\pm$0.6
Sokal 93.0$\pm$0.6 72.3$\pm$3.4 33.9$\pm$1.2 12.4$\pm$0.5 5.6$\pm$0.6
Tanimoto 93.1$\pm$0.6 72.8$\pm$1.0 35.1$\pm$1.2 14.0$\pm$0.5 7.5$\pm$0.6
No communication Cosine 95.2$\pm$1.1 79.2$\pm$1.3 44.2$\pm$1.2 20.6$\pm$0.7 12.2$\pm$0.9
Dice 95.1$\pm$1.1 79.1$\pm$1.3 44.0$\pm$1.2 20.4$\pm$0.7 11.8$\pm$0.9
Russel 94.8$\pm$1.2 77.6$\pm$1.4 40.2$\pm$1.2 14.5$\pm$0.9 4.2$\pm$1.1
Sokal 94.9$\pm$1.1 78.0$\pm$1.4 41.2$\pm$1.2 16.0$\pm$0.8 6.2$\pm$1.1
Tanimoto 95.0$\pm$1.1 78.4$\pm$1.3 42.2$\pm$1.2 17.6$\pm$0.8 8.2$\pm$1.0
Appendix E Additional Discussion about Isomers
E.1 Isomer Category
@@
Isomers typically fall into two main categories: constitutional (structural) isomers, which share the same chemical formula but display distinct atom connectivity, and stereoisomers (spatial isomers), which share the same topology graph but diverge in their three-dimensional arrangement (see Figure E.1). Constitutional isomers are NMR-variant, meaning that different isomers produce distinct NMR spectrum. In the sub-categories of stereoisomers, enantiomers are NMR-invariant, but diastereomers and cis-trans isomers are NMR-variant.
Figure E.1: NMR Variability in Isomers
E.2 Isomers Group for $\text{C}_{7}\text{H}_{11}\text{NO}_{3}$
@@
Here is an example for isomer groups. In this isomer group of $\text{C}_{7}\text{H}_{11}\text{NO}_{3}$, they all share the same chemical formula in Figure E.2. The first 10 are constitutional (structural) isomers of each other (cycled green), the last 4 are two pairs of diastereomers (cycled brown). Each of these isomers corresponds to a distinct NMR spectrum.
Figure E.2: Isomer demo for $\text{C}_{7}\text{H}_{11}\text{NO}_{3}$
Appendix F Additional Results on Peak Assignment
@@
In complex natural product molecules, it is a common situation that the local contents of some atoms within the same molecule exhibit a high degree of similarity. It gives rise to challenges for the atomic alignment, as some atoms correspond to ppm values in close proximity. However, our K-M3AID model is capable of recognizing each of the atoms with effective learnt embeddings and deciphering the correspondences among the atoms and the peaks at zero-shot. Two complex natural product molecules with multiple rings (4 and 4, respectively) and multiple chiral centers (6 and 8, respectively) are taken to showcase the effectiveness of atomic alignment (see Appendix Figure F.1).
@@
Figure F.1: Examples of Zero-shot Atomic Alignment for Complex Natural Products. Yellow cells in the PPM difference represent the ground truth alignment.
In molecular A in Figure  F.2, atom 13 and atom 14 are tertiary carbons (attaching to 3 carbons and 1 hydrogen) and on the same 5-member ring, corresponding to the ppm of 34.3 and 35.6, respectively. The similar local content of these two atoms fools SP-ID and WP-ID. In addition, WP-ID fails with more atomic alignments. The molecular B is chemical symmetric regarding atom 0. Thus, atom 1 and atom 3 correspond to the same peak on the spectra. The ppm of atom 1 and atom 3 is 114.2, the ppm of atom 2 and atom 4 is 110.0. While there is 4.2 difference, SP-ID and WP-ID fails to pick up right alignment for atom 1 and atom 3. In contrast, K-ID succeed to align the atoms with peaks in both molecules.
@@
Figure F.2: Extra case studies of IE-Meta-MMA.Yellow cells in the PPM differerence represent the ground truth alignment, and red cross represents the wrong alignment.